20191016_ DX ESM science

06:58 Send DX ESM Preset

07:06 Loop closed, LUCI2 collect science
  • 15.3 mag
  • 100 Hz
  • bin 4
  • manual gains 1.5/0.75/0.75

07:10 DIMM seeing 1.2 and variable

07:09 and beyond telemetry collected
  • 20191016_*
  • 070934 Loop closed, pause, loop closes 4000 frames
  • 071047 Loop closed, pause 4000 frames
  • 071136 All closed 2000 frames
  • 071305 All closed 2000 frames
  • 071617 OK
  • 071953 OK

07:20 My impression from the WFS and residual RMS is the correct is reasonable and stable (should be about the same all the time, even thought the seeing is changing).

07:21 Program done. Loop opened

-- DougMiller - 16 Oct 2019
Topic revision: r1 - 16 Oct 2019, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback